OpenStudy (anonymous):

Help Please?? I'm confused. Input in standard form the equation of the given line. The line that passes through (-3, 5) and (-2, -6)

OpenStudy (jack1):

y = mx + c is standard form? or is that gradient intercept form?

OpenStudy (anonymous):

for mine it says the standard form y=mx+b

OpenStudy (jack1):

ok, then points are on the cartesian plane are always written x then y \[ \large so~ (-3, 5) = (x_1, y_1)~ and~ (-2, -6) = (x_2, y_2) \] therefore y = mx + c \[\large m = \frac {y_2 - y_1}{x_2 - x_1} \]

OpenStudy (jack1):

and b (or c, whichever = y intercept) find that by subbing in x 1 and y1 y = m(x) + c -6 = m(-2) + c sub in ur above value for m ...only c is left

OpenStudy (anonymous):

Okay, i got it from here, thank you got your help!